OpenStudy (m.d.saxon):

some people go to a party there is a total of 5 handshakes. How many people are at the party?

OpenStudy (kainui):

"Party" "5 handshakes"

OpenStudy (kainui):

Try drawing out a picture with dots, and lines between the dots represent handshakes.

OpenStudy (anonymous):

Dots representing handshakes? You just blew my mind.

OpenStudy (kainui):

no no no, lines are the handshakes. The dots are the people.

OpenStudy (anonymous):

dats wut i ment

OpenStudy (anonymous):

sorry ma english is not so gud

OpenStudy (kainui):

I don't know how looking at how many considerate people are attending a party has anything to do with the total number of people attending.

OpenStudy (anonymous):

It doesn't lol, I think op made this question up or is missing something.

OpenStudy (mathmale):

Here's my point of view: Only two people are involved in any one handshake. Let's say that there are n people at the party; we don't know what n is yet. All we know is that five separate handshakes take place. Think "combinations." \[_{n}C _{x}\]represents the number of combinations possible from a set of n, taken x at a time. It happens that this is equal to \[\frac{ n! }{ x!(n-x)! }\] So here, with n unknown, but x known to be 2, and the total number of combos known to be 5, we write the equation\[\frac{ n! }{ 2!(n-2)! }=5\] and solve it for the single unknown, n. Anyone?

OpenStudy (mathmale):

It occurs to me that n has to be at least 2, because of that (n-2)! term in the denominator.

OpenStudy (mathmale):

Note: I've tried n = 2, 3, 4 and 5, but so far have not found that any of these possible values for n "works." Comments? Suggestions?

OpenStudy (triciaal):

4 with one missing handshake